40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 1211     Entry time: Wed Feb 25 04:29:49 2015
Author: Zach 
Type: DailyProgress 
Category: SiFi - ringdown 
Subject: Taiwan cantilever has higher Q, going for cryo cycle now 

[Nic, Zach]

Nic got a Glasgow-style cantilever from a group in Taiwan, and a quick test in the rapid cycle chamber showed that it had pretty low loss, so we are running it in the cryostat now. As a reminder, these are the rough dimensions of this style cantilever:

Below is a photo of the box it came in, showing the actual 92-um thickness of this sample, as well as a shot of it in the vacuum chamber. For some reason, this particular sample's clamping tab did not fit in the groove that Nic had built into the clamping block for the other Glasgow cantilevers, so I had to mount it to the side against the flat faces of the clamp (as I've been doing with our larger samples).

 

This evening, we transferred it over to the cryostat and restored all the electrical connections for what will hopefully be a fruitful cryo run. Here is a ringdown of the fundamental mode (~106 Hz) at room temperature:

The measured decay time of 41 seconds corresponds to a Q of around 14,000, which is about as good as we expect at room temperature. This sample is probably better than our other ones for at least 2 reasons:

  1. It is made from a better-quality (FZ) wafer, and
  2. It has been manufactured monolithically with a thicker clamping tab, which our modeling suggests is a very effective way to evade clamping losses by keeping strain energy within the silicon.

Given that we didn't see much improvement at all with our other samples when going to low temperature, I believe (2) is by far the biggest effect. The Glasgow wafers only have the clamp-region thickness extended to one side, which is modelled to be worse than if you go both ways, but it is still much better than we can do with our discrete sandwiching.

I filled the LN2 reservoir and the volume is cooling overnight. I did some rough ringdowns at a point when the steel block was registering around 160 K and found greatly improved Qs already (approaching and perhaps exceeding 105). We will continue to make measurements tomorrow.

ELOG V3.1.3-